a 40 m x 20 m swimming pool is filled with water t a depth of 2.8m, determine the heat in kj required to raise the temperature of the water in the pool from 12degree C to 29 degree C


1
Expert's answer
2021-11-11T17:08:29-0500

Total volume of water "= 40 \times 20 \times 2.8 = 2240 \;m^2"

Density of water ρ= 1000 kg/m3

Mass of water "m = 2240 \times 1000 = 2.24 \times 10^6 \; kg"

Specific heat of water "s = 4184\; J/kg \; K"

ΔT = 29-12 = 17 °C

The required amount of heat

"Q = msΔT \\

Q = 2.24 \times 10^6 \times 4184 \times 17 = 159326 \times 10^6 \;J \\

Q= 159326 \;MJ"